OpenStudy (anonymous):

Steve and Alice realized that their ages are consecutive odd integers. The product of their ages is 195. Steve is younger than Alice. Determine their ages by completing the square.

zepdrix (zepdrix):

Woops! Caleb has the right idea, but 195 is a PRODUCT of the 2 numbers. So it would be x(x+1)=195

OpenStudy (anonymous):

I made a oopsy :(

zepdrix (zepdrix):

hehe, it happens ^^

OpenStudy (anonymous):

Ugh I dont get it =/

zepdrix (zepdrix):

Let Steve's age be X. If their ages are CONSECUTIVE, then Alice's age is the number that comes directly after X. So it will be X+1. Example: If steve is 15, alice is 16 .. or to put it another way, alice is 15+1. This is what they tell us in the problem: (steve's age)*(alice's age)=195 So let's plug in the variables we chose for steve and alice's age. (X)*(X+1)=195 Understand where this equation is coming from now? :O

OpenStudy (anonymous):

thanks!

zepdrix (zepdrix):

Do you know how to solve for X from that spot? :) Requires a few algebra steps.

OpenStudy (anonymous):

complete the square right?

zepdrix (zepdrix):

yah c: multiply the stuff out, complete square stuff. fun times.

OpenStudy (anonymous):

Please check me after I do it :)

OpenStudy (anonymous):

I solve this right just to be sure x^2+x-195=0?

zepdrix (zepdrix):

Hmmmmm yah :o completing the square.. hmm my brain is rusty for some reason lol

OpenStudy (anonymous):

lol its okay Ima do it real quick correct me if im wrong.

zepdrix (zepdrix):

OH CRAP sorry I led you astray D:::

zepdrix (zepdrix):

Now I see why the numbers are not working out.

zepdrix (zepdrix):

Their ages are consecutive ODD integers.

zepdrix (zepdrix):

So if steve is 15, alice is 15+2, or 17. So if steve is X, alice is X+2. That's the equation we want to start with X(X+2)=195 OK ok ok ok that will work out a lot better :) yay sorry about that.

OpenStudy (anonymous):

thank you!